# Single-Image based unsupervised joint segmentation and denoising

### v1

@@ 5 Experiments and Results
In Figure 10, we show another example of image segmentation for three different noise levels using Algorithm 4. The main difficulty of this image lies in the intensities which are shared by the object to be segmented and the background. Therefore, we chose a representative box for initialising the squirrel, which includes both, dark and bright areas, in order to enable the foreground denoising expert to better generalize on the foreground region consisting of dark and bright areas. Naturally, as the squirrel and background do not differ much in terms of their structural properties, the foreground denoiser, $\Phi_{\theta^{F}}$ also performs well on the background, causing the segmentation mask $u$ to grow. In order to control this behaviour, we applied our second strategy that includes a recursive reference mask as described in Algorithm 3, thus preventing the segmentation mask obtained at iteration $k+1$ from deviating too much from the previous one at iteration $k$. More precisely, the parameters that we used for noise level 10 were $\mu=0.0001,\lambda=0.005$, for noise level 30 we set $\mu=0.005$, $\lambda=0.005$, while for a noise level 50 $\mu=0.00015$, $\lambda=0.005$.
Figure 10: Segmentation results obtained on the noisy images showing a squirrel corrupted with three different noise levels. The first column shows clean input image, and initialisation for foreground and background regions, while in the second column the noisy versions of the given image are depicted. The remaining ones present the segmentation results obtained using the proposed strategy with the segmentation Algorithm 4, the segmentation masks using the Chan-Vese algorithm provided by skimage with checkerboard initialisation and box initialisation, respectively. The last row shows the denoised images which are obtained by composing the obtained segmentation mask and expert denoiser outputs.
@@ 6 Extensions and limitations
First, our proposed unified framework can be extended to the (multichannel) multiclass segmentation case, as we discuss in the following paragraph.
@@ 6.1 Vector-valued multi-class model
In order to segment a noise-corrupted vector-valued image represented as $\boldsymbol{f}=(f_{1},\dots,f_{L})$ into $C$ different regions, we can consider $C$ dedicated neural networks acting as denoising experts for each region. In this case, the objective is to estimate $C$ segmentation masks $\{u_{i}\}_{i=1}^{C}$ satisfying the simplex constraint, i.e. $\sum_{k=1}^{C}u_{i}=1$, as well as the set of network parameters $\boldsymbol{\theta}^{\text{MC}}=(\theta^{\text{MC}}_{1},\dots,\theta^{\text{MC}}_{C})$. With these notations, the energy (5) can be extended to segment noise-corrupted, vector-valued images $\boldsymbol{f}$ as
$\mathcal{E}_{f,\lambda}(\boldsymbol{u},\boldsymbol{\theta})\coloneqq i_{\mathbb{A}}(\boldsymbol{u})$ $+\lambda\lvert\boldsymbol{u}\rvert_{\text{TV}}+\sum_{i=1}^{C}\sum_{j=1}^{L}\int_{\Omega}\left(f_{j}-\Phi_{\theta_{i}^{\text{MC}}}(f_{j})\right)^{2}u_{i}\,.$ (13)
@@ 6.2 Limitations
A limitation of the current work lies in the training strategy in the case where two denoisers are applied. In our experiments, we observed that once the denoising experts have been trained on the initial boxes and the subsequent segmentation step has been realised, it may occur that one of the two classes in the image contains an important part of the other class. As a result, during the next denoising step, one of the network is trained on parts of both regions present in the image. With the influence of the total variation regularisation, $u$ may converge to an undesired constant mask. With the recursive integration of a reference mask, we already proposed in section 4.2 a strategy to overcome this drawback. One interesting alternative would be to include an additional constraint enforcing the denoisers to perform better in their initial regions than in the other initial ones.
Next, in some of our experiments we have observed that the Noise2Fast denoiser is not well suited for the segmentation of certain images such as the zebra image in Figure 3. The reason for that is that the filters of the learned experts operate locally, and are not good in capturing global information of the regions. As a consequence, in the case of the zebra, in regions where the stripes are thicker, and the intensity values are closer to the one in the background, the background expert outperforms the one of the foreground, resulting in an undesired result as obtained by the piecewise constant Chan-Vese model. To overcome this limitation, we modified the checkerboard strategy of the Noise2Fast method, and instead of squeezing the image to half of its width/height, we divided its size by a factor of four. In addition to include different denoisers, such as for instance the deep image prior [39], an interesting perspective would be to define new data fitting terms focusing on structural similarities within the different classes.

### v2

@@ 5 Experiments and Results
In Figure 10, we show another example of image segmentation for three different noise levels using Algorithm 4. The main difficulty of this image lies in the intensities which are shared by the object to be segmented and the background. Therefore, we chose a representative box for initialising the squirrel, which includes both, dark and bright areas, in order to enable the foreground denoising expert to better generalize on the foreground region consisting of dark and bright areas. Naturally, as the squirrel and background do not differ much in terms of their structural properties, the foreground denoiser, $\mathbf{D}_{F}(f)$ also performs well on the background, causing the segmentation mask $u$ to grow. In order to control this behaviour, we applied our second strategy that includes a recursive reference mask as described in Algorithm 3, thus preventing the segmentation mask obtained at iteration $k+1$ from deviating too much from the previous one at iteration $k$. More precisely, the parameters that we used for noise level 10 were $\mu=0.0001,\lambda=0.005$, for noise level 30 we set $\mu=0.005$, $\lambda=0.005$, while for a noise level 50, we considered $\mu=0.00015$ and $\lambda=0.005$.
Figure 10: Segmentation results obtained on the noisy images showing a squirrel corrupted with three different noise levels. The first column shows clean input image, and initialisation for foreground and background regions, while in the second column the noisy versions of the given image are depicted. The remaining ones present the segmentation results obtained using the proposed strategy with the segmentation Algorithm 4, the segmentation masks using the Chan-Vese algorithm provided by skimage [33] with checkerboard initialisation and box initialisation, respectively. The last column shows the denoised images which are obtained by composing the obtained segmentation mask and expert denoiser outputs.
@@ 6 Extensions and limitations
In this scetion, we first show how our unified framework can be extended to the (multichannel) multiclass segmentation case. Then we discuss the current limitations of the proposed joint denoising and segmentation approach.
@@ 6.1 Vector-valued multi-class model
In order to segment a noise-corrupted vector-valued image represented as $\boldsymbol{f}=(f_{1},\dots,f_{L})$ into $C$ different regions, we can consider $C$ dedicated neural networks acting as denoising experts for each region. In this case, the objective is to estimate $C$ segmentation masks $\{u_{i}\}_{i=1}^{C}$ satisfying the simplex constraint, i.e. $\sum_{k=1}^{C}u_{i}=1$, as well as the set of network parameters $\boldsymbol{W}^{\text{MC}}=(W^{\text{MC}}_{1},\dots,W^{\text{MC}}_{C})$. With these notations, the energy (5) can be extended to segment noise-corrupted, vector-valued images $\boldsymbol{f}$ as
$\mathcal{E}_{f,\lambda}(\boldsymbol{u},\boldsymbol{W})\coloneqq i_{\mathbb{A}}(\boldsymbol{u})$ $+\lambda\lvert\boldsymbol{u}\rvert_{\text{TV}}+\sum_{i=1}^{C}\sum_{j=1}^{L}\int_{\Omega}\left(f_{j}-\mathbf{D}_{W_{i}^{\text{MC}}}(f_{j})\right)^{2}u_{i}\,.$ (13)
@@ 6.2 Limitations
One limitation is encountered when training two denoisers, where one might encompass parts of both regions during segmentation, leading to undesired mask convergence. Although we propose a strategy to address this, adding a constraint to prioritize denoising performance in initial regions could be beneficial.
Moreover, the Noise2Fast denoiser may not be suitable for certain images, such as the zebra example, due to its local filter operation. This can result in outcomes similar to the piecewise constant Chan-Vese model. We modified the Noise2Fast strategy and suggest exploring alternative denoisers, such as the deep image prior [32]. Additionally, using contrastive learning ideas, introducing new data fitting terms focusing on structural similarities within classes could enhance segmentation performance.